OpenStudy (anonymous):

HELP PLEASEEEE x y 2 1 3 1 4 -3 3 -3 Part A: Does the table represent a relation that is a function? Justify your answer by using the values shown in the table. Part B: The function f(h) shown below represents Kylie's total earning from working in a cafe for h hours of work. f(h) = 15h + 5 Calculate and interpret the meaning of f(8)? Part C: Write an ordered pair to represent the input and output of the function in Part B when Kylie works for r hours.

OpenStudy (anonymous):

Part A: Yes X goes in Y comes out.

OpenStudy (anonymous):

F(h) just means Y

OpenStudy (anonymous):

Yeah im still very clueless...

OpenStudy (anonymous):

No since X= 3 has two images

OpenStudy (anonymous):

What ?

OpenStudy (anonymous):

I literally do not understand any of this.

OpenStudy (anonymous):

@Jvital it's not a function

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

I know its not... can we move on from part a please

OpenStudy (anonymous):

f(8)=15.8+5 she receive 15$ per hour+ an initial amount

OpenStudy (anonymous):

thank you mr.smarty pants, now help me with part c please

OpenStudy (anonymous):



OpenStudy (anonymous):

is that the answer for part c ? (r,15r+5) ? it looks strange..

OpenStudy (anonymous):

hello ?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

how did you get that odd answer ?

OpenStudy (anonymous):

oh never mind,i see now... kind of.

OpenStudy (anonymous):

@RaphaelFilgueiras can you help me with another question ?

OpenStudy (anonymous):

i will try

OpenStudy (anonymous):

Ok, here it is The table below represents a linear function f(x) and the equation represents a function g(x). x f(x) -1 -5 0 -1 1 3 g(x) g(x) = 4x+3 Part A: Write a sentence to compare the slope of the two functions and show the steps you used to determine the slope of f(x) and g(x) Part B: Which function has a greater y-intercept? Justify your answer

OpenStudy (anonymous):

@TheBadWolf here it is

OpenStudy (anonymous):

@TheBadWolf @RaphaelFilgueiras hello ? boys ?

OpenStudy (anonymous):

To find the slope we need to use the slope formula\[\frac{ y_{2}-y _{1} }{ x _{1}-x _{2} }\] and f(x) is basically just y so we can just input values from the chart so\[\frac{ -5 - (-1) }{ -1-0 }\]

OpenStudy (anonymous):

i thought we established i have no idea how to do any of this /: i get the slope but it says to compare and how the hell do you find the slope of a stupid chart

OpenStudy (anonymous):

That would equal to -4/-1 which is just 4 so thats the slope for f(x) The slope for g(x) is easy because it is right infront of the x in the equation which is 4. so the slopes are the same making the lines parallel

OpenStudy (anonymous):

to find the intercepet plug in x=0

OpenStudy (anonymous):

That is all for part a. For part b, g(x) has a greater y-intercept because when the x values for each eqaution is 0 f(x)=-1 as given by the chart and g(x)=3 which can be seen in the equation

OpenStudy (anonymous):

ok this is what im putting as my final answer part a : the slopes are the same meaning they form parallel lines. −5−(−1) ------- = -4/-1 = 4 -1-0 slope of g(x)= 4 part b : g(x) has a greater y-intercept because when the x values for each eqaution is 0 f(x)=-1 as given by the chart and g(x)=3 which can be seen in the equation

OpenStudy (anonymous):

i basically copied and pasted because i really cannot stand algebra anymore.

OpenStudy (anonymous):

@TheBadWolf last one, i promise ! Part A: Sam rented a boat at $225 for 2 days. If he rents the same boat for 5 days, he has to pay a total rent of $480. Write an equation in the standard form to represent the total rent (y) that Sam has to pay for renting the boat for x days. (4 points) Part B: Write the equation obtained in Part A using function notation.(2 points) Part C: Describe the steps to graph the equation obtained above on the coordinate axes. Mention the labels on the axes and the intervals. (4 points)

OpenStudy (anonymous):

I could swear I did this question already

OpenStudy (anonymous):

or at least saw it

OpenStudy (anonymous):

y=225-2x y=480-5x and solved for y got y=55 That's the first part

OpenStudy (anonymous):

I think

OpenStudy (anonymous):

I'm not really sure this one is confusing

OpenStudy (anonymous):

yeah i submitted it already, im so done with this class, i need help with oneeee more though :P ill tag you in it